# A First Look at Information Highlighting in Stack Overflow Answers

### v1

@@ 4.1 RQ1: How prevalent is the information highlighting in SO answers?
Approach: To understand the prevalence of information highlighting in SO answers, we compute the percentage of answers that have the studied formatting types (see Table 1) and the percentage of words highlighted in each answer with each formatting. In addition, we compute the basic statistics of information highlighting instances to understand the characteristics of each formatting type. More specifically, we calculate the distribution of instances for each formatting and the number of words highlighted with them.
@@
Bold
11.3%
2.0/1.0/241
9.0%/4.1%/100%
12.0%
2.9/1.0/436
Italic
7.2%
1.9/1.0/354
5.9%/2.3%/100%
7.3%
2.9/1.0/477
Delete
0.07%
1.2/1.0/36
18.1%/10.6%/100%
0.04%
15.7/8.0/701
Code
38.5%
3.7/2.0/490
8.7%/6.3%/100%
78.9%
1.4/1.0/4,669
Heading
1.6%
2.1/2.0/168
9.3%/3.9%/100%
1.7%
3.4/2.0/327
@@ 4.2 RQ2: What types of information are highlighted with Code formatting in SO answers?
Approach: In this RQ, we aim to understand what content is highlighted with the Code. First, we randomly sampled 385 Code instances with a 5% interval and 95% confidence level. Since there is no existing terminologies we could reuse for this purpose, we manually performed a lightweight open coding-like process Seaman (1999); Zhang et al. (2019) to identify the type of content highlighted with Code. The process involves three phases and is performed by the first authors (A1 and A2) of the paper. In phase I, A1 and A2 derived a draft list of types based on 50 Code instances. During the phase, the types were revised and refined. In phase II, A1 and A2 independently applied the derived types to the rest 335 samples. They took notes regarding the diffidence or ambiguity during the labeling. During this phase, no new types were introduced. In phase III, A1 and A2 discussed the results from Phase II to resolve any disagreements until a consensus was reached. The coding process has a Cohen’s kappa of 0.8 (measured before starting Phase III), which indicates a substantial level of the inter-rater agreement Viera et al. (2005). Table 3 presents the definition and the corresponding example of the derived types for Code.
@@
Identifier
The identifier in source code, e.g, “ArrayList”.
245 (63.6%)
Keyword
The keywords in programming languages, e.g., “for” and “public”.
38 (9.9%)
Statement
A statement of source code, e.g., “plot = last.plot()”.
27 (7.0%)
Equation
Mathematical equation, operator, or number, e.g., “O(log n)”.
21 (5.2%)
Software
The name of softwares, frameworks, tools, and libs, e.g., “Tensorflow” and “MySQL”.
19 (4.9%)
Path
Path and files name, e.g., “src/main/java”.
17 (4.4%)
Terminology
Terminologies related to programming, e.g., “hash table”.
7 (1.8%)
Cmd
Command, e.g., “cloud-init init”
5 (1.3%)
Version
Version information of a software, e.g., “62.1”.
2 (0.5%)
Other
Other than the above defined types.
4 (1%)
@@
Approach: In this RQ, we aim to understand what content is highlighted with the studied Text formatting (i.e., Bold, Italic, Delete, and Heading). Similar to RQ2, we performed an manual study. We randomly sampled 385 Text instances with a 5% interval and 95% confidence level. We ended up with 177 Bold, 169 Italic, 3 Delete, and 36 Heading instances. We then identified the types of highlighted content by following the methodology used in RQ2. The coding process has a Cohen’s kappa of 0.78 (measured before starting Phase III), which indicates a substantial level of the inter-rater agreement Viera et al. (2005). Table 4 presents the definition and the corresponding example of the derived types for Text formatting.
@@
Update
The update and new edit in the post, e.g., “Update: According to your comments, I think you will want to try gzfile() in read.table()”.
22 (12.4%) 2 (1.2%) 0 0
Equation
Same as the definition of Equation in Table 3.
3 (1.7%) 2 (1.2%) 0 0
Source code
The union of Identifier/Keyword/Statement/Operator/Cmd/Path in Table 3, e.g., “I’d suggest the usage of startupOrder to configure the startup of route though”.
51 (28.8%) 52 (30.8%) 0 1 (2.8%)
Caveat
A reminder or warn of in which context or condition the provided solution works or does not work, e.g., “It won’t work in that shell though, you need to open a new one”.
33 (18.6%) 43 (25.4%) 0 1 (2.8%)
Reference
Reference to internal or external resource, e.g., “<a href="http://api.jquery.com/not/">.not()</a>”.
12 (6.8%) 17 (10.1%) 0 0
Results
Results or output.
3 (1.7%) 2 (1.2%) 0 0
Terminology
Same as the definition of Terminology in Table 3.
12 (6.8%) 21 (12.4%) 0 0
Heading
The title of a section, e.g., “WORKING EXAMPLE”.
27 (15.3%) 2 (1.2%) 0 33 (91.6%)
Options
Options in software, e.g., “you have to enable Less Secure Sign-In in your google account”.
4 (2.3%) 1 (0.6%) 0 0
Extent
A word/phrase to express extent, e.g., “at least with Hibernate it assumes you want to use a global "hibernate" sequence for all tables, which is just stupid.”
4 (2.3%) 14 (8.3%) 0 0
Version
Same as the definition of Version in Table 3.
0 3 (1.8%) 0 0
Delete
Deleted outdated/wrong description, e.g., “Just use KVO KVC”.
0 0 3 (100%) 0
Other
Other than the above defined types.
6 (3.4%) 10 (5.9%) 0 1
Results: Both Bold and Italic formatting are most frequently used to highlight content related to source code. Table 4 presents the distribution of each Text type. We observe for certain types, Bold and Italic share similar patterns. For instance, in 28.8% of the Bold instances and 30.8% of the Italic instances, content related source code (i.e., Source code) is highlighted. That is said, users also frequently use Bold and Italic to highlight source code other than using Code formatting. Users probably use Bold, Italic, and Code exchangeably, although SO suggests users to tag inline code using Code format StackExchange (2023). Interestingly, we also observe in some cases, the community members changed formatting of the code-related content from Bold and Italic to Code. For instance, in an answer, a user change the formatting for a funciton “strncmp()” from Italic to Code [^7] https://stackoverflow.com/posts/18437465/revisions.
@@ Results
Users tend to highlight the content of types Update and Heading with Bold, while Italic is more likely to be used for the type of Extent. From Table 4, we can see differences between Bold and Italic. For instance, Bold is more frequently used to highlight Update (12.4%) and Heading (15.3%) compare with Italic. Interestingly, users also use Bold to highlight a heading. One possible reason is that the rendering effect for Bold and Heading are similar. Prior studies reveal that answers on SO are easy to become obsolete Ragkhitwetsagul et al. (2019); Zhang et al. (2019). Therefore, it is not surprising to observe that users highlight the update using Text formatting. Different from Bold, Italic is used more frequently to highlight a word/phrase to express the extent (i.e., Extent).
@@ 4.4 RQ4: Can we recommend the highlighted content in SO answers automatically?
Approach: In the previous RQs, we focused on analyzing the highlighted content on Stack Overflow. As the highlighting is prevalent and the information type can be categorized, some suggestions based on these uses could benefit the user of Stack Overflow. The objective is to construct a recommendation system designed to propose potential words to users, along with suitable tags, with the aim of enhancing the user’s reading experience and comprehension. In this specific RQ, we explore the feasibility of implementing an automated machine learning-based method for content highlighting. This feature would be helpful, particularly for SO users who are inexperienced in the practice of highlighting information and seek to automatically highlight critical content when composing their posts. Specifically, we aim to create a model that can highlight information in an input answer post using appropriate formatting types (Bold, Italic, Delete, Code, and Heading). To achieve this, we use neural network architectures initially designed for Named Entity Recognition (NER) without actually training NER models. Our task shares similarities with NER, such as the short length of labeled (highlighted) tokens and a limited number of text labeling (formatting) options. Moreover, both NER and automated information highlighting can be formed as a sequence labeling task. Note that we train the neural networks from scratch on our information highlighting dataset.
Figure 2: Layers of the NER model
In this RQ, we have used two different structured models for comparison. One model is based on the Convolutional Neural Network (CNN) model provided by spaCySPACY (2023) and another is a transformer model, BERT Devlin et al. (2019).
Figure 2 presents the architecture of our CNN model. In this paper, we use the Token2Vec model SpaCy (2023) provided by a well-known open-source software library for advanced Natural Language Processing (NLP) tasks named spaCy SPACY (2023). We opt to utilize spaCy since it has emerged as the de facto standard for practical NLP due to its speed, robustness, and performance Le Guillarme and Thuiller (2022). The model architecture comprises a token embedding layer (128 dimensions), four convolutional layers, and an output layer. The token embedding layer is named as MultiHashEmbed by spaCy. It features an embedding layer that embeds a number of lexical attributes separately. Then the results are concatenated as a vector semantic representation of a token for passing through a feed-forward network. As it uses hashing to store embeddings in memory, it also supports faster processing of large datasets. The embedding layer is followed by a 4-layered Convolutional Neural Network (CNN), named MaxoutWindowEncoder. Each of the convolutional layers is followed by a max pooling layer. The final layer is a softmax layer for predicting labels on tokens. Training employs cross-entropy loss and the Adam optimizer.
Figure 3: Layers of the NER model by Bert
On the other hand, BERT is a state-of-the-art transformer model pre-trained on a large dataset Devlin et al. (2019). It can be fine-tuned for numerous natural language processing tasks, encompassing question answering, language inference, text classification, and sequence tagging. The architecture of the model is based on a multi-layer bidirectional Transformer encoder developed by researchers of Google. The model considers the contexts of both preceding and subsequent words. The architecture of BERT consists of multiple transformer blocks, each containing a self-attention mechanism for getting the context of the input sequence. In this work, we fine-tuned BERT for our prediction task. We used the pre-trained model from HuggingFace platform Face (2023).
Experiment Setup: We construct the dataset for this RQ by selecting posts that contained highlighted words from Section 3. Then we divided the post into sentences and extracted the corresponding formatting tags described in Section 3. In other words, each data point is a highlighted instance (a sentence containing at least one type of information highlighting). When preparing the training data, we consider the content highlighted with different formatting types as the entities in the NER model. We annotate the tagged words with their formatting tags, considering their position in the phrase. For example, in the sample highlighted sentence shown in Figure 2, i.e., “You will have to use client_wait_for().”, “client_wait_for()” is formatted with a code tag. To label this sentence, we assign the label “B-code” to the highlighted word, indicating that it is the beginning of a highlighted Code content, while the other words in the sentence are labeled as O (Outside). If multiple words are formatted, we used I (Interior) and E (Ending) to label the highlighted content’s internal words and the ending word. For example, “you need to run.git push --force --all first” contains more than one word in the highlighted Code content. So the labels for the phrase would be (B-code: “git”), (I-code: “push”), (I-code: “--force”) and (E-code: “--all”). Note that we train separate models for each highlighting type. When preparing the dataset for a particular highlighting type, if a sentence has other highlighting types rather than the one we target, we omit other types of highlighting types from the sentence. Table 5 presents the basic statistics of our dataset for different formatting types.
Table 5: Number of sentences in train and test datasets.
Type Training set Testing set
Bold
2,590,193
646,486
Italic
1,669,344
417,409
Code
9,292,312
2,324,836
Heading
406,946
102,133
Following prior studies, we partitioned the dataset into an 80:20 ratio for training and testing purposes. We trained our models in 3 epochs where the learning rate is 0.001. We employ two training settings.
We use 32 as the size of the batch in the fixed-size setting (i.e., fixed). Models based on BERT architecture are trained in 0.05 learning rate using 16 batches(fixed) in 6 epochs.
Table 6: Models parameters.
Parameter CNN BERT
Epochs
3
6
Learning rate
0.001
0.05
Batch size
32
16
Evaluation: We use precision, recall, and F1-score to measure the performance of models, following prior studies evaluating NER models Esuli and Sebastiani (2010); Jiang et al. (2016). Instead of using the exact match (i.e., a true positive is counted when all of the tokens of a named entity are labeled correctly, both the length and type), we decide to use a partial match Nadeau and Sekine (2007); Tanabe et al. (2005), in which we consider if each individual token is predicted correctly. The partial match allows us to measure the performance of the segmentation aspects of highlighted content when the content has multiple words. More specifically, we calculate the precision and recall as follows:
@@
Results: Out of the four trained models, the Code model obtained the highest F1 score of 0.71 (spaCy). Table 7 presents the results of our trained models in terms of precision, recall, and F1 using CNN and BERT sizes of batches. In general, in terms of F1, we observe that the Code models (F1 score is 0.343 and 0.71 using BERT and CNN models) perform better than other models. This is reasonable, since first for code formatting, we have much more training data than other types. Secondly, code typically has certain special characters (e.g., “.”, and “()” in API “class.function()”) and keywords (e.g., “foreach”), making it easy for the model to identify. Our observation indicates that it is easier to build a model to identify code and highlight them on SO posts than other types of formatting.
On the other hand, BERT archives higher precision(0.699 to 0.92) than spaCy models however the recalls are below 0.05 which is the lowest recall in spaCy models. With this low recall, it ends with the F1 scores lower than the spaCy models. Only the BERT model that was trained for Code performs scores 0.343 in F1 Scoring, which is also lower than the code models of spaCy. So it can be said that spaCy models performed better than BERT models.
@@
Table 7: The performance of models on all formatting types Bold, Italic, Heading, Code with different training settings. F denotes the fixed models and I denotes the incremental models.
@@
$Model_{Bold}$
0.76
0.76
0.11
0.025
0.19
0.049
$Model_{Italic}$
0.82
0.92
0.05
0.003
0.10
0.006
$Model_{Heading}$
0.62
0.84
0.21
0.0093
0.31
0.018
$Model_{Code}$
0.78
0.699
0.65
0.228
0.71
0.343
Our CNN-based models achieve an acceptable precision ranging from 0.62 to 0.82 for different formatting types while obtaining low recall. Overall, except for the Code models, other models have higher precision and lower recall. For models for text formatting, Heading, Bold, and Italic models can achieve at least a precision of 0.62 from the models. While models for text formatting struggle from low recall. The results suggest that as long as the text formatting models recommend the content to highlight, the accuracy is acceptable, while the models miss most of the content that is supposed to be highlighted (low recall).
@@
Bold
NOTE : THIS MIGHT ALREADY SOMETHING THAT YOU HAVE TRIED. This is probably due to the fact that you are using –trusted-host=pypi.python.org.
NOTE : THIS MIGHT ALREADY SOMETHING THAT YOU HAVE TRIED. This is probably due to the fact that you are using –trusted-host=pypi.python.org.
Code
For more info on searching for the elements on DOM such as getElementById,querySelector, please refer here.
For more info on searching for the elements on DOM such as getElementById,querySelector, please refer here.
Italic
Side note: Your code is falling prey to what I call The Horror of Implicit Globals because you never declare filename.
Side note: Your code is falling prey to what I call The Horror of Implicit Globals because you never declare filename.
Heading
OPTION 1: GIT STASH
You may postpone all current changes of your files in the git stash cache. After that the files are equal to former checkout state.
OPTION 1: GIT STASH
You may postpone all current changes of your files in the git stash cache. After that the files are equal to former checkout state.
@@
    Misidentification: Although the content that needs highlight is successfully identified, incorrect formatting types are recommended. In other words, the content is not identified by the correct model.
@@
    False Identification: The model identifies the content that is not supposed to highlight.
@@
    Missing Identification: The model misses the content that is supposed to highlight.
@@
Table 10 presents the distribution of those three types of failures across different models. Overall, the majority of the failure cases are missing identification (e.g., 55% for incremental model) and false identification (e.g., 39% for incremental model), while misidentification takes the least portion (e.g., 6.2% for incremental model). All models suffer from high missing identification except Code models, which explains the low recall in Table 7, typically for Bold and Italic models. 89% and 66% of the failures in Bold and Italic CNN models are missing identification. One possible explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge) Kandpal et al. (2022); Mireshghallah et al. (2022). To test this assumption, we count the frequency of the words that are correctly predicted and the words that are not identified by models in the training data. Table 9 presents the mean/median frequency of words that are correctly predicted and those that are not identified by models. On top of it, Figure 4 compares their distributions. We observe that the mean and median frequency of words that are predicted correctly is much larger than those of words that are not identified in all models, except the median value of Code.
@@
Bold
4.05/1.0
193.05/22.0
Italic
6.11/1.0
674.48/130.0
Heading
2.53/1.0
34.73/5.0
Code
1.35/1.0
5.37/1.0
@@
Failure Type Formatting Type
Bold
Code
Italic
Heading
Total
Misidentification
6.8%
6.7%
1.6%
12%
6.2%
Missing Identification
89%
39%
66%
59%
55%
False Identification
4%
54%
32%
28%
39%
Code, Italic, and Heading models tend to have false identification failures, especially the Code models. For instance, in example 7 of Table 11, the Code model identify “setTimeout” while it is not highlighted in the original text. setTimeout is indeed a function name and one possible explanation is that setTimeout is frequently highlighted with Code in the training data, and the model can recognize it. Similarly, in example 4, the Italic model identifies “I currently” after “Disclaimer:” as Italic. Recall example 3 in Table 8, in the original text, “Disclaimer: I work at SerpApi” was italicized. The model may identify the word “Disclaimer:” and tend to italicize the content after it.
@@
We also observe a small portion of misidentification cases. For instance, in example 1 in Table 11, the original text is highlighted with Heading, while our model identifies the content but formats it using bold. One possible explanation is that those types of formatting share common usage patterns to some extent and confuse the models. We actually observed such a phenomenon in Section 4.3. For instance, we observe that both Bold and Italic are used to highlight source code and Caveat. Heading is used to highlight source code.
@@
Table 11: Failure cases of spacy detection of tags for all the types Bold, Italic, Heading, and code.
ID Type Original text
@@
Text highlighted with our models (Incremental).
Failure type
1
Bold
Swift 5.2.x
@@
First of all, you need to declare an "easy to use" typealias for your block:
Swift 5.2.x
First of all, you need to declare an "easy to use" typealias for your block:
Misidentification: Detected heading as bold.
2
Bold
@@
If you want to extend it’s expiration date, IT WILL THROW THE EXCEPTION!
If you want to extend it’s expiration date, IT WILL THROW THE EXCEPTION!
False Identification: detected normal texts as bold.
3
Bold
@@
From now, your frontend application will use access token in the Authorization header for every request.
From now, your frontend application will use access token in the Authorization header for every request.
Missing Identification: missed Bold.
4
Italic
Disclaimer: I currently work for Mapbox.
Disclaimer: I currently work for Mapbox.
@@
False Identification: detected normal texts as Italic.
5
Heading
@@
Close Sublime and Quit It. Open up your desired project that you want to use Sublime with in Unity
Close Sublime and Quit It. Open up your desired project that you want to use Sublime with in Unity
False Identification: detected normal texts as Heading.
6
Heading
@@
Because there are no spaces in the text. Chinese is commonly written without any space characters between words or sentences. The ASS subtitle format, having been originally designed by and for European language speakers, unfortunately only breaks words on actual space characters. Your subtitle file does not contain any space characters, so the lines are never broken.
Because there are no spaces in the text. Chinese is commonly written without any space characters between words or sentences. The ASS subtitle format, having been originally designed by and for European language speakers, unfortunately only breaks words on actual space characters. Your subtitle file does not contain any space characters, so the lines are never broken.
Missing Identification: Missing Heading for “Because there are no spaces in the text”.
7
Code
@@
setTimeout seems like the right guy for the job. To off a click use jQuery’s .off() Method. Also, the .one() method to attach a callback only once :
setTimeout seems like the right guy for the job. To off a click use jQuery’s .off() Method. Also, the .one() method to attach a callback only once :
False Identification: Detected normal text as code.
8
Code

### v3

@@ 5.1 RQ1: How prevalent is the information highlighting in SO answers?
Bold 11.3% 2.0/1.0/241 9.0%/4.1%/100% 12.0% 2.9/1.0/436
Italic 7.2% 1.9/1.0/354 5.9%/2.3%/100% 7.3% 2.9/1.0/477
Delete 0.07% 1.2/1.0/36 18.1%/10.6%/100% 0.04% 15.7/8.0/701
Code 38.5% 3.7/2.0/490 8.7%/6.3%/100% 78.9% 1.4/1.0/4,669
Heading 1.6% 2.1/2.0/168 9.3%/3.9%/100% 1.7% 3.4/2.0/327
@@ 5.2 RQ2: What types of information are highlighted with Code formatting in SO answers?
Identifier The identifier in source code, e.g, “ArrayList”. 245 (63.6%)
Keyword The keywords in programming languages, e.g., “for” and “public”. 38 (9.9%)
Statement A statement of source code, e.g., “plot = last.plot()”. 27 (7.0%)
Equation Mathematical equation, operator, or number, e.g., “O(log n)”. 21 (5.2%)
Software The name of softwares, frameworks, tools, and libs, e.g., “Tensorflow” and “MySQL”. 19 (4.9%)
Path Path and files name, e.g., “src/main/java”. 17 (4.4%)
Terminology Terminologies related to programming, e.g., “hash table”. 7 (1.8%)
Cmd Command, e.g., “cloud-init init” 5 (1.3%)
Version Version information of a software, e.g., “62.1”. 2 (0.5%)
Other Other than the above defined types. 4 (1%)
@@
Update The update and new edit in the post, e.g., “Update: According to your comments, I think you will want to try gzfile() in read.table()”. 22 (12.4%) 2 (1.2%) 0 0
Equation Same as the definition of Equation in Table 5. 3 (1.7%) 2 (1.2%) 0 0
Source code The union of Identifier/Keyword/Statement/Operator/Cmd/Path in Table 5, e.g., “I’d suggest the usage of startupOrder to configure the startup of route though”. 51 (28.8%) 52 (30.8%) 0 1 (2.8%)
Caveat A reminder or warn of in which context or condition the provided solution works or does not work, e.g., “It won’t work in that shell though, you need to open a new one”. 33 (18.6%) 43 (25.4%) 0 1 (2.8%)
Reference Reference to internal or external resource, e.g., “<a href="http://api.jquery.com/not/">.not()</a>”. 12 (6.8%) 17 (10.1%) 0 0
Results Results or output. 3 (1.7%) 2 (1.2%) 0 0
Terminology Same as the definition of Terminology in Table 5. 12 (6.8%) 21 (12.4%) 0 0
Heading The title of a section, e.g., “WORKING EXAMPLE”. 27 (15.3%) 2 (1.2%) 0 33 (91.6%)
Options Options in software, e.g., “you have to enable Less Secure Sign-In in your google account”. 4 (2.3%) 1 (0.6%) 0 0
Extent A word/phrase to express extent, e.g., “at least with Hibernate it assumes you want to use a global "hibernate" sequence for all tables, which is just stupid.” 4 (2.3%) 14 (8.3%) 0 0
Version Same as the definition of Version in Table 5. 0 3 (1.8%) 0 0
Delete Deleted outdated/wrong description, e.g., “Just use KVO KVC”. 0 0 3 (100%) 0
Other Other than the above defined types. 6 (3.4%) 10 (5.9%) 0 1
Results: Both Bold and Italic formatting are most frequently used to highlight content related to source code. Table 6 presents the distribution of each Text type. We observe for certain types, Bold and Italic share similar patterns. For instance, in 28.8% of the Bold instances and 30.8% of the Italic instances, content related source code (i.e., Source code) is highlighted. That is said, users also frequently use Bold and Italic to highlight source code other than using Code formatting. Users probably use Bold, Italic, and Code exchangeably, although SO suggests users to tag inline code using Code format [41]. Interestingly, we also observe in some cases, the community members changed formatting of the code-related content from Bold and Italic to Code. For instance, in an answer, a user change the formatting for a funciton “strncmp()” from Italic to Code [^8] https://stackoverflow.com/posts/18437465/revisions.
@@
Users tend to highlight the content of types Update and Heading with Bold, while Italic is more likely to be used for the type of Extent. From Table 6, we can see differences between Bold and Italic. For instance, Bold is more frequently used to highlight Update (12.4%) and Heading (15.3%) compare with Italic. Interestingly, users also use Bold to highlight a heading. One possible reason is that the rendering effect for Bold and Heading are similar. Prior studies reveal that answers on SO are easy to become obsolete [33, 55]. Therefore, it is not surprising to observe that users highlight the update using Text formatting. Different from Bold, Italic is used more frequently to highlight a word/phrase to express the extent (i.e., Extent).
@@ 5.4 RQ4: Can we recommend the highlighted content in SO answers automatically?
Out of the four trained models, the Code model obtained the highest F1 score of 0.69 (CNN). Table 7 presents the results of our trained models in terms of precision, recall, and F1 using CNN and BERT sizes of batches. In general, in terms of F1, we observe that the Code models (F1 score is 0.162 and 0.69 using BERT and CNN models) perform better than other models. This is reasonable, since first for code formatting, we have much more training data than other types. Secondly, code typically has certain special characters (e.g., “.”, and “()” in API “class.function()”) and keywords (e.g., “foreach”), making it easy for the model to identify. Our observation indicates that it is easier to build a model to identify code and highlight them on SO posts than other types of formatting.
We also notice that for code data, CNN-based models outperform BERT-based models by a large margin, typically in terms of recall. The CNN-based model for code achieves a recall of 0.66, while the BERT-based model only has a recall of 0.093. In other words, the CNN-based model recognizes and covers more real code-related content than the BERT-based model. One possible explanation is that the CNN-based model is smaller than BERT and makes it easier to train and learn the patterns to recognize the code-related content to highlight. Another possibility is that BERT is pre-trained on natural language and not for code, which makes it hard to recognize code-related content.
@@
Table 7: The performance of models on all formatting types Bold, Italic, Heading, Code with different training settings. CNN denotes the CNN-based models and BERT denotes the BERT-based models.
@@
$Model_{Bold}$ 0.57 0.63 0.05 0.006 0.09 0.011
$Model_{Italic}$ 0.64 0.53 0.01 0.0025 0.02 0.005
$Model_{Heading}$ 0.50 0.49 0.04 0.0073 0.08 0.0143
$Model_{Code}$ 0.72 0.65 0.66 0.093 0.69 0.162
Our CNN-based models achieve an acceptable precision ranging from 0.50 to 0.72 for different formatting types while obtaining low recall. Overall, except for the Code models, other models have higher precision and lower recall. For models for text formatting, Bold, and Italic models can achieve at least a precision of 0.57 from the models. While models for text formatting struggle from low recall. The results suggest that as long as the text formatting models recommend the content to highlight, the accuracy is acceptable, while the models miss most of the content that is supposed to be highlighted (low recall). One possible reason is the training data for those three types are much smaller than code and the patterns for text formatting types are more diverse, which makes it challenging to learn models for text formatting.
@@
Bold NOTE : THIS MIGHT ALREADY SOMETHING THAT YOU HAVE TRIED. This is probably due to the fact that you are using –trusted-host=pypi.python.org. NOTE : THIS MIGHT ALREADY SOMETHING THAT YOU HAVE TRIED. This is probably due to the fact that you are using –trusted-host=pypi.python.org.
Code For more info on searching for the elements on DOM such as getElementById,querySelector, please refer here. For more info on searching for the elements on DOM such as getElementById,querySelector, please refer here.
Italic Side note: Your code is falling prey to what I call The Horror of Implicit Globals because you never declare filename. Side note: Your code is falling prey to what I call The Horror of Implicit Globals because you never declare filename.
Heading OPTION 1: GIT STASH You may postpone all current changes of your files in the git stash cache. After that the files are equal to former checkout state. OPTION 1: GIT STASH You may postpone all current changes of your files in the git stash cache. After that the files are equal to former checkout state.
@@
    Misidentification: Although the content that needs highlighting is successfully identified, incorrect formatting types are recommended. In other words, the content is not identified by the correct model.
@@
    False Identification: The model identifies the content that is not supposed to be highlighted.
@@
    Missing Identification: The model misses the content that is supposed to be highlighted.
@@
Failure Type Formatting Type
Bold Code Italic Heading
Misidentification 0.86% 4.09% 2.65% 15.19%
Missing Identification 95.6% 52.77% 96.67% 81.1%
False Identification 3.55% 43.13% 0.67% 3.63%
Table 9 presents the distribution of those three types of failures across different models. Overall, the majority of the failure cases are missing identification All models suffer from high missing identification except Code models, which explains the low recall in Table 7, typically for text formatting tag (Bold, Italic, and Heading) models. One possible explanation is that the models are easy to learn the frequent highlighted words while struggling to learn less frequent words (i.g., long-tail knowledge) [18, 24]. To test this assumption, we count the frequency of the words that are correctly predicted and the words that models in the training data do not identify. We examine the recommendation results produced by each CNN-based model. Table 10 presents the mean/median frequency of words that are correctly predicted and those that are not identified by models. On top of it, Figure 6 compares their distributions. We observe that the mean and median frequency of words that are predicted correctly is much larger than those of words that are not identified in all models, except the median value of Code.
@@
Bold 127.7/7 441.5/37.0
Italic 2.39/2.52 959.9/332.0
Heading 1.86/1.94 241.24/89.0
Code 50% 0.699/0.602 1931.04/8.0
@@
We also observe a small portion of misidentification cases and False identification. For instance, in example 1 in Table 11, the original text is highlighted with Heading, while our model identifies the content but formats it using bold. One possible explanation is that those types of formatting share common usage patterns to some extent and confuse the models. We actually observed such a phenomenon in Section 5.3. Users sometimes highlight the same content with both Bold Italic tags. For instance, there are many cases in which if there is any update on the posts, the keyword “update” will be highlighted by different types of formatting, i.e., “update” was highlighted the most by Bold tag (0.028%) followed by other tags (Heading 0.0054%, Code 0.00269% and Italic 0.00249%). Therefore, when recommending the information to highlight, if multiple models recommend the same content (e.g., both the Bold model and Italic model recommend “update” to be highlighted), we use the formatting tag that is returned by the model with the highest confidence as the final tag. This decision is based on the assumption that the recommendation from the model with the highest confidence aligns with the predominant choice of the community. By prioritizing the recommendation from the most confident model, we aim to ensure consistency and accuracy in the formatting of highlighted information, reflecting the collective preferences of the user community. Another option is that we can return all recommendation results with confidence scores returned by each model and allow users to select, like Google Query Suggestion [1].
@@
Table 11: Failure cases predicted by CNN-based models for all the types Bold, Italic, Heading, and code.
@@
ID Type Original text Text highlighted with our models (Incremental). Failure type
@@
1 Bold Swift 5.2.x First of all, you need to declare an "easy to use" typealias for your block: Swift 5.2.x First of all, you need to declare an "easy to use" typealias for your block: Misidentification: Detected heading as bold.
@@
2 Bold If you want to extend it’s expiration date, IT WILL THROW THE EXCEPTION! If you want to extend it’s expiration date, IT WILL THROW THE EXCEPTION! False Identification: detected normal texts as bold.
@@
3 Bold From now, your frontend application will use access token in the Authorization header for every request. From now, your frontend application will use access token in the Authorization header for every request. Missing Identification: missed Bold.
@@
4 Italic Disclaimer: I currently work for Mapbox. Disclaimer: I currently work for Mapbox. False Identification: detected normal texts as Italic.
@@
5 Heading Close Sublime and Quit It. Open up your desired project that you want to use Sublime with in Unity Close Sublime and Quit It. Open up your desired project that you want to use Sublime with in Unity False Identification: detected normal texts as Heading.
@@
6 Heading Because there are no spaces in the text. Chinese is commonly written without any space characters between words or sentences. The ASS subtitle format, having been originally designed by and for European language speakers, unfortunately only breaks words on actual space characters. Your subtitle file does not contain any space characters, so the lines are never broken. Because there are no spaces in the text. Chinese is commonly written without any space characters between words or sentences. The ASS subtitle format, having been originally designed by and for European language speakers, unfortunately only breaks words on actual space characters. Your subtitle file does not contain any space characters, so the lines are never broken. Missing Identification: Missing Heading for “Because there are no spaces in the text”.
@@
7 Code setTimeout seems like the right guy for the job. To off a click use jQuery’s .off() Method. Also, the .one() method to attach a callback only once : setTimeout seems like the right guy for the job. To off a click use jQuery’s .off() Method. Also, the .one() method to attach a callback only once : False Identification: Detected normal text as code.